# Hidden Metallic Iron in Amorphous Silicate Dust? Insights from Condensation Experiments and Mid-Infrared Spectroscopy

### v1

@@ 5 Conclusions
Condensation experiments using the ITP system were performed in the Mg–Ca–Na–Al–Si–Fe–Ni–O and Mg–Ca–Na–Al–Si–O systems to synthesize two types of amorphous silicate dust analogues with Solar composition: amorphous silicate with metallic Fe-Ni core, which has a distribution of core size proportion from 0 to 0.87, and Fe-Ni-free amorphous silicate.
Reflectance spectra were measured in addition to absorbance spectra to derive the optical constants. The powders were pressed into pellets, and reference materials were prepared by coating the surface of sample pellets with gold. This approach effectively prevented a decrease in reflectance due to the surface roughness. We derived optical constants from the reflectance at ¿4000 cm-1 to determine and fitting calculated from absorbance spectra.
Using the determined optical constants, the effects of metallic iron particles in amorphous silicate grains on IR spectral features were discussed. The modeled spectra of core-shell dust showed clear shifts of the Si–O stretching and Si–O–Si bending peaks with larger metallic cores (). However, the absorbance spectra of experimental products exhibited no difference in the peak positions regardless of the presence or absence of metallic cores. This is because Fe and Ni were concentrated in metallic cores () in large amorphous silicate grains () and small amorphous silicate grains () with small cores () or without cores dominate the spectral features. As a result, the presence of metallic iron particles cannot be detected from infrared spectra. The spectral effects of metallic iron depend more on the size distribution and number density of inclusions than their elemental abundance. Metallic Fe-Ni grains formed on the surface of silicate melts sink into the silicate melts and coagulate into a large core to save surface energy, resulting in a size distribution of metallic cores.
The metallic-iron-bearing dust analogue produced in this study is comparable to GEMS in terms of iron abundance and the distribution of metallic core sizes. However, even in the presence of abundant iron, the mid-infrared spectra are indistinguishable from those of pure amorphous silicate dust when the metallic cores have a size distribution. This implies that dust around AGB stars may not be composed solely of pure amorphous silicates, but could also contain metallic iron cores similar to those in GEMS. The combination of MIR and NIR spectroscopic observations and radiative transfer modeling using the optical constants obtained in this study will constrain the presence of circumstellar GEMS-like materials.
@@ Acknowledgements.
H. Enomoto and A. Takigawa were financially supported by JSPS KAKENHI Grant Number 23H01213. H. Enomoto was supported by JST SPRING, Grant Number JPMJSP2108. The authors thank Dr. Ryo Tazaki of The University of Tokyo for constructive discussion of the spectral modeling of dust with metallic core.Appendix A Optical constants of the metallic iron-free product CI-2 (Ca0.05Mg0.92Na0.06Al0.07Si)

### v2

@@ 5 Conclusions
We performed condensation experiments using the ITP system in the Mg–Ca–Na–Al–Si–Fe–Ni–O and Mg–Ca–Na–Al–Si–O systems to synthesize two types of amorphous silicate dust analogs with a solar composition: amorphous silicate with a metallic Fe-Ni core, which has a core size proportion () from 0 to 0.87, and Fe-Ni-free amorphous silicate. Reflectance spectra were measured in addition to absorbance spectra to derive the optical constants. The powders were pressed into pellets, and the reference materials were prepared by coating the surface of the sample pellets with gold. This approach effectively prevented the decrease in reflectance that would have been caused by the surface roughness. We derived optical constants by determining from the reflectance at cm-1 and fitting the calculated from the model to that derived from the measured absorbance.
Using the determined optical constants, the effects of metallic iron particles in amorphous silicate grains on IR spectral features were discussed. The modeled spectra of core-shell dust showed clear shifts of the Si–O stretching and Si–O–Si bending peaks with larger metallic cores (). However, the peak positions of the absorbance spectra of the experimental products remained the same regardless of the presence or absence of metallic cores. This is because Fe and Ni were concentrated in metallic cores () in large amorphous silicate grains () and because small amorphous silicate grains () with small cores () or without cores dominate the spectral features. As a result, the presence of metallic iron particles cannot be detected from IR spectra. The spectral effects of metallic iron depend more on the size distribution and number density of inclusions than on their elemental abundance. Metallic Fe-Ni grains formed on the surface of silicate melts sink into the silicate melts and coagulate into a large core to minimize surface energy, thereby leading to a size distribution of metallic cores.
The metallic-iron-bearing dust analog produced in this study is comparable to GEMS in terms of iron abundance and the distribution of metallic core sizes. However, even in the presence of abundant iron, the MIR spectra are indistinguishable from those of pure amorphous silicate dust when the metallic cores have a size distribution. This implies that the dust around AGB stars is not composed solely of pure amorphous silicates but could also contain metallic iron cores similar to those in GEMS. Combining radiative transfer modeling using the optical constants obtained in this study with MIR and NIR spectroscopic observations provides a powerful means to constrain the presence of circumstellar GEMS-like materials.
@@ Acknowledgements.
H. Enomoto and A. Takigawa were financially supported by JSPS KAKENHI Grant Number 23H01213. H. Enomoto was supported by JST SPRING, Grant Number JPMJSP2108. The authors thank Dr. Ryo Tazaki (The University of Tokyo) for constructive discussions on spectral modeling of dust with metallic cores. We also thank an anonymous referee for their insightful comments that improved the quality of this paper.Appendix A Optical constants of the metallic iron-free product CI-2 (Ca0.05Mg0.92Na0.06Al0.07Si)